# Weak convergence of Bayes estimators under general loss functions

### v1

@@ 5 Examples
We review here several popular loss functions that behave locally like a polynomial and thus fall within our framework. A natural and widely used class is given by $\ell(t,\vartheta)=\left\lVert t-\vartheta\right\rVert^{p},\;p>0,$ which clearly satisfies the assumptions of Theorem 2. Special cases include the posterior mean for $p=2$ and the posterior (geometric) median for $p=1$. The asymptotic behavior of Bayes estimators under such losses has been extensively studied, and in fact, stronger results are known. For instance, Chao, (1970) showed that the difference between a Bayes estimator corresponding to these loss functions and the maximum likelihood estimator converges to zero at a rate faster than $n^{-1/2}$. This directly implies that the two estimators share the same asymptotic distribution.
@@ 5.1 Locally quadratic loss functions
In contrast, for an intrinsic loss $\ell$, the corresponding Bayes estimator is parametrization invariant, as shown in Robert, (1996, Lemma 6.2). Formally, let $\mathcal{M}=\{P_{\vartheta}:\vartheta\in\Theta\}$ be a fixed parametric family of probability measures, each absolutely continuous with respect to Lebesgue measure on $\mathbb{R}^{d}$. The parametrization map $\Psi:\Theta\to\mathcal{M}$ is given by $\vartheta\mapsto P_{\vartheta}$. We denote by $p_{t}$ the Lebesgue density of $P_{t}$. A few examples of intrinsic losses that satisfy the assumptions of Theorem 3 for some suitable parametric family are:
@@
    and $\mathcal{P}(\mathcal{X}\times\mathcal{X})$ is the set of all probability measures on $\mathcal{X}\times\mathcal{X}$ with finite second moment;
  3. (C)
    Parametric Kullback–Leibler divergence:
    $\ell_{\mathrm{KL}}(t,\vartheta)=\mathrm{KL}\bigl(\Psi(\vartheta)\mid\Psi(t)\bigr)=\int_{\mathcal{X}}\log\!\left(\frac{p_{\vartheta}(x)}{p_{t}(x)}\right)\mathop{}\!\mathrm{d}P_{\vartheta}(x).$
It is shown in Robert, (1996, Lemma 4.1) that (A) and (C) behave locally like a quadratic function.
2-Wasserstein distance. Since the behavior of intrinsic loss functions depends on the underlying parametric family, we give an example of a parametric family for which $\ell_{W_{2}}$ in (B) satisfies the conditions of Theorem 3 beyond the location–scale setting, in which case the loss function reduces to the Euclidean distance.
@@
and thus the loss function has a well-separated minimum, which together with the triangle inequality implies that Assumption 2 holds. Thus, the conditions of Theorem 3 are satisfied.
More general differentiability conditions for loss functions induced by the 2-Wasserstein distance, together with concrete examples, are given in Section A.3.
Sinkhorn divergence. Entropy-regularized Wasserstein distances are widely used due to their computational efficiency (see e.g. Peyré and Cuturi,, 2019). We give a concrete example to show that such loss functions also fit within our framework. For a regularization parameter $\lambda>0$ and a parametric family $\{P_{\vartheta}:\vartheta\in\Theta\}$ on $\mathcal{X}$ with finite second moments, define the Sinkhorn divergence $S_{\lambda}$ with squared Euclidean cost between $P_{t}$ and $P_{\vartheta}$ as
@@
We consider the loss function $\ell(t,\vartheta):=d_{S,\lambda}\!\left(P_{t},P_{\vartheta}\right)$. If the model is identifiable (i.e., $P_{t}=P_{\vartheta}$ implies $t=\vartheta$), then $\ell(t,\vartheta)\geq 0$ with equality if and only if $t=\vartheta$ (Lavenant et al.,, 2024).
Let $\mathcal{A}=\{(\sigma_{1},\sigma_{2},\sigma_{3})\in(0,\infty)^{3}:\sigma_{1}\sigma_{3}>\sigma_{2}^{2}\}$, and consider the bivariate Gaussian family
@@
From this, the loss function $\ell((\mu,\sigma),(\nu,\tau))$ can be expressed explicitly, and one can see that it is quadratic in $\mu$ and thrice continuously differentiable in $\sigma$, with a unique minimum at $(\mu,\sigma)=(\nu,\tau)$. Therefore, $\ell$ satisfies the assumptions of Theorem 3.
@@ 5.2 Stein discrepancies
Total variation distance. A special case of an integral probability metric in (6) is the total variation distance, obtained with $\mathcal{H}=\{\mathbb{I}_{A}:A\in\mathcal{B}(\mathcal{X})\}$. In Section 4, we considered loss functions satisfying a local quadratic expansion and showed that condition (iii) in Proposition 3 holds for a quadratic loss $\ell_{0}$. We now present an example of a loss with a different local behavior, where the corresponding $\ell_{0}$ is not quadratic.
Example 3.
Let $\{P_{\vartheta}:\vartheta\in\Theta\subseteq\mathbb{R}^{d}\}$ be a parametric family with densities $p_{\vartheta}=\mathop{}\!\mathrm{d}P_{\vartheta}/\mathop{}\!\mathrm{d}\mu$ with respect to a $\sigma$-finite measure $\mu$ on $\mathcal{X}$. Assume the family is $L^{1}(\mu)$-differentiable, i.e., there exists $\dot{p}_{\vartheta}:\mathcal{X}\to\mathbb{R}^{d}$ such that $\left\lVert p_{\vartheta+r}-p_{\vartheta}-\langle r,\dot{p}_{\vartheta}\rangle\right\rVert_{L^{1}}=\int_{\mathcal{X}}\left\lvert p_{\vartheta+r}(x)-p_{\vartheta}(x)-\langle r,\dot{p}_{\vartheta}(x)\rangle\right\rvert\mathop{}\!\mathrm{d}\mu(x)=o(\left\lVert r\right\rVert),$ as $r\to 0$ in $\mathbb{R}^{d}$. For the loss $\ell(t,\vartheta)=2\left\lVert P_{t}-P_{\vartheta}\right\rVert_{\mathsf{TV}}=\left\lVert p_{t}-p_{\vartheta}\right\rVert_{L^{1}}$, we then have $\ell(t,\vartheta)=\left\lVert\langle t-\vartheta,\dot{p}_{\vartheta}\rangle\right\rVert_{L^{1}}+R(t,\vartheta),$ where $R(t,\vartheta)=o(\left\lVert t-\vartheta\right\rVert)$ as $\left\lVert t-\vartheta\right\rVert\to 0$.
Differentiability in quadratic mean, assumed in the BvM theorem (Theorem 1), implies $L^{1}$-differentiability, and moreover, $\dot{p}_{\vartheta}(x)=p_{\vartheta}(x)\ell^{\prime}_{\vartheta}(x)$, where $\ell^{\prime}_{\vartheta}$ is the quadratic mean derivative. Assume that $\vartheta\mapsto\dot{p}_{\vartheta}(x)$ is continuous at $\vartheta_{0}$ for all $x\in\mathcal{X}$, that for some $\delta>0$
$\int_{\mathcal{X}}\sup_{\vartheta\in B_{\delta}(\vartheta_{0})}\left\lVert\dot{p}_{\vartheta}(x)\right\rVert\,\mathop{}\!\mathrm{d}\mu(x)<\infty,$
and that for any compact $K\subset\mathbb{R}^{d}$ and $h\in\mathbb{R}^{d}$,
@@
Then condition (iii) in Proposition 3 holds with $\ell_{0}(t,h)=\left\lVert\langle t-h,\dot{p}_{\vartheta_{0}}\rangle\right\rVert_{L^{1}}\text{ and }p=1.$ The reverse triangle inequality and the Cauchy–Schwarz inequality yield uniform convergence as follows
which converges to zero, as $\varepsilon\to 0$, by continuity. Thus, if the other assumptions of Theorem 2 hold, the Bayes estimator under total variation loss satisfies
$n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})\stackrel{{\scriptstyle D}}{{\to}}\arg\min_{t\in\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left\lVert\langle t-h,\dot{p}_{\vartheta_{0}}\rangle\right\rVert_{L^{1}}\,\mathop{}\!\mathrm{d}\mathcal{N}\!\left(Y,I_{\vartheta_{0}}^{-1}\right)(h),$
@@
A direct verification (in Section C.5), using $L^{1}$-differentiability and uniform directional differentiability of the Gaussian density, shows that for any compact $K\subset\mathbb{R}^{d}$ and $h\in\mathbb{R}^{d}$,
@@
1-Wasserstein distance. For the particular choice of Stein class $\mathcal{H}=\{f:\mathcal{X}\to\mathbb{R}:\left\lvert f(x)-f(y)\right\rvert\leq\left\lVert x-y\right\rVert\text{ for any }x,y\}$ with compact $\mathcal{X}$, the induced integral probability metric in (6) is the $1$-Wasserstein distance (Villani,, 2003, Theorem 1.14). If $\mathcal{X}$ is one-dimensional (not necessarily compact), the distance admits the distribution function representation (Villani,, 2003, Theorem 2.18): For $P,Q$ with finite first absolute moments and distribution functions $F_{P},F_{Q}$, it holds that $W_{1}(P,Q)=\int_{\mathbb{R}}\left\lvert F_{P}(x)-F_{Q}(x)\right\rvert\,\mathop{}\!\mathrm{d}x=\left\lVert F_{P}-F_{Q}\right\rVert_{L^{1}}.$
@@
If the family is differentiable in quadratic mean, then under mild conditions $F_{\vartheta}$ is $L^{1}$-differentiable with $\dot{F}_{\vartheta}(t)=\int_{-\infty}^{t}\ell^{\prime}_{\vartheta}(x)\,\mathop{}\!\mathrm{d}P_{\vartheta}(x),$ where $\ell^{\prime}_{\vartheta}$ is the quadratic mean derivative (Proposition 6).
@@
Figure 1: Contour plots of mappings $(a,b)\mapsto W_{1}(P_{a},P_{b})$ (left) and $(a,b)\mapsto\left\lvert a-b\right\rvert$ (right).
@@ 5.3 Beyond our framework
Here, we cannot employ our framework, since the empirical prior $\Pi_{n}$ is not absolutely continuous with respect to the Lebesgue measure (so that the BvM theorem may fail). Nevertheless, this barycenter can be formulated as the solution of a finite-dimensional linear program, whose standard form (see Section C.5) is
@@
The barycenter is then expressed as $t=(x_{1},\ldots,x_{d-1},\,1-x_{1}-\cdots-x_{d-1})$. Its limiting distribution (which is possibly non-Gaussian) may be studied via arguments analogous to those in Klatt et al., (2022), who provided limit laws for linear programs with random $b$ in contrast to the random $c$ encountered here. A detailed investigation of this problem is an intriguing direction for future research, but lies beyond the scope of this paper.
@@ 6 Numerical experiments
In this section we illustrate our theoretical convergence results through numerical simulations. We focus on the three intrinsic losses $\ell_{H},\ell_{W_{2}},\ell_{\mathrm{KL}}$ discussed in Section 5.1.
@@ 6.1 Exponential–Gamma model
$\ell_{H}(t,\vartheta)=1-\frac{2(t\vartheta)^{1/2}}{t+\vartheta},\quad\ell_{W_{2}}(t,\vartheta)=2(t^{-1}-\vartheta^{-1})^{2},\quad\ell_{\mathrm{KL}}(t,\vartheta)=\log(\vartheta)-\log(t)+\frac{t}{\vartheta}-1.$
In the simulation, we set $a=b=2$ and $\vartheta_{0}=2$. Figure 2 displays QQ-plots of the rescaled and centered estimators $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ against the standard normal distribution. As $n$ increases, the sample quantiles approach the theoretical normal quantiles, in agreement with Theorem 3.
Figure 2: $Q$-plots for $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ for $n\in\{10,10^{2},10^{4},10^{6}\}$ (left to right) based on $M=500$ Monte Carlo repetitions, where $\vartheta_{0}=2$ and $\hat{\theta}_{n}$ is the Bayes estimator under losses $\ell_{H}$, $\ell_{W_{2}}$ and $\ell_{\mathrm{KL}}$ (top to bottom). The reference corresponds to the standard normal distribution.

### v2

@@ 5 Examples
We review here several popular loss functions that behave locally like a polynomial and thus fall within our framework. A natural and widely used class is given by $\ell(t,\vartheta)=\left\lVert t-\vartheta\right\rVert^{p},\;p>0,$ which clearly satisfies the assumptions of Theorem 2. Special cases include the posterior mean for $p=2$ and the posterior (geometric) median for $p=1$. The asymptotic behavior of Bayes estimators under such losses has been extensively studied, and in fact, stronger results are known. For instance, as shown in Chao, (1970), the difference between a Bayes estimator corresponding to these loss functions and the maximum likelihood estimator converges to zero at a rate faster than $n^{-1/2}$. This directly implies that the two estimators share the same asymptotic distribution.
@@ 5.1.1 Intrinsic loss functions
In contrast, for an intrinsic loss $\ell$, the corresponding Bayes estimator is parametrization invariant, as shown in Robert, (1996, Lemma 6.2). Formally, let $\mathcal{M}=\{P_{\vartheta}:\vartheta\in\Theta\}$ be a fixed parametric family of probability measures, each absolutely continuous with respect to Lebesgue measure on $\mathbb{R}^{d}$. The parametrization mapping $\Psi:\Theta\to\mathcal{M}$ is given by $\vartheta\mapsto P_{\vartheta}$. We denote by $p_{t}$ the Lebesgue density of $P_{t}$. A few examples of intrinsic losses that satisfy the assumptions of Theorem 3 for some suitable parametric family are:
@@
    Parametric Kullback–Leibler divergence:
    $\ell_{\mathrm{KL}}(t,\vartheta)=\mathrm{KL}\bigl(\Psi(\vartheta)\mid\Psi(t)\bigr)=\int_{\mathcal{X}}\log\!\left(\frac{p_{\vartheta}(x)}{p_{t}(x)}\right)\mathop{}\!\mathrm{d}P_{\vartheta}(x);$
  3. (C)
@@
    and $\mathcal{P}(\mathcal{X}\times\mathcal{X})$ is the set of all probability measures on $\mathcal{X}\times\mathcal{X}$ with finite second moment.
It is shown in Robert, (1996, Lemma 4.1) that (A) and (B) behave locally like a quadratic function.
5.1.2 2-Wasserstein distance
Since the behavior of intrinsic loss functions depends on the underlying parametric family, we give an example of a parametric family for which $\ell_{W_{2}}$ in (C) satisfies the conditions of Theorem 3 beyond the location–scale setting, in which case the loss function reduces to the Euclidean distance.
@@
and thus the loss function has a well-separated minimum, which together with the triangle inequality implies Assumption 2. Thus, the conditions of Theorem 3 are satisfied.
More general differentiability conditions for loss functions induced by the 2-Wasserstein distance, together with concrete examples, are given in Supplement A.3.
5.1.3 Sinkhorn divergence
Entropy-regularized Wasserstein distances are widely used due to their computational efficiency (see e.g. Peyré and Cuturi,, 2019). We give a concrete example to show that such loss functions also fit within our framework. For a regularization parameter $\lambda>0$ and a parametric family $\{P_{\vartheta}:\vartheta\in\Theta\}$ on $\mathcal{X}$ with finite second moments, define the Sinkhorn divergence $S_{\lambda}$ with squared Euclidean cost between $P_{t}$ and $P_{\vartheta}$ as
@@ 5.1.3 Sinkhorn divergence
We consider the loss function $\ell(t,\vartheta):=d_{S,\lambda}\!\left(P_{t},P_{\vartheta}\right)$. If the model is identifiable (i.e., $P_{t}=P_{\vartheta}$ implies $t=\vartheta$), then $\ell(t,\vartheta)\geq 0$ with equality if and only if $t=\vartheta$, see Lavenant et al., (2024).
Let $\mathcal{A}=\{(\sigma_{1},\sigma_{2},\sigma_{3})\in(0,\infty)^{3}:\sigma_{1}\sigma_{3}>\sigma_{2}^{2}\}$, and define the bivariate normal family
@@
From this, the loss $\ell((\mu,\sigma),(\nu,\tau))$ can be expressed explicitly, and one can see that it is quadratic in $\mu$ and thrice continuously differentiable in $\sigma$, with a unique minimum at $(\mu,\sigma)=(\nu,\tau)$. Therefore, the loss $\ell$ satisfies the assumptions of Theorem 3.
@@ 5.2 Stein discrepancies
5.2.1 Total variation distance
A special case of an integral probability metric in (7) is the total variation distance, obtained with $\mathcal{H}=\{\mathbb{I}_{A}:A\in\mathcal{B}(\mathcal{X})\}$. In Section 4, we considered loss functions satisfying a local quadratic expansion and showed that condition (iii) in Proposition 3 holds for a quadratic loss $\ell_{0}$. We now present an example of a loss with a different local behavior, where the corresponding $\ell_{0}$ is not quadratic.
Example 3 ($L^{1}$-differentiable family).
Let $\{P_{\vartheta}:\vartheta\in\Theta\subseteq\mathbb{R}^{d}\}$ be a parametric family with densities $p_{\vartheta}=\mathop{}\!\mathrm{d}P_{\vartheta}/\mathop{}\!\mathrm{d}\mu$ with respect to a $\sigma$-finite measure $\mu$ on $\mathcal{X}$. Assume that the family is $L^{1}(\mu)$-differentiable, i.e. there exists $\dot{p}_{\vartheta}:\mathcal{X}\to\mathbb{R}^{d}$ such that
$\left\lVert p_{\vartheta+r}-p_{\vartheta}-\langle r,\dot{p}_{\vartheta}\rangle\right\rVert_{L^{1}}=\int_{\mathcal{X}}\left\lvert p_{\vartheta+r}(x)-p_{\vartheta}(x)-\langle r,\dot{p}_{\vartheta}(x)\rangle\right\rvert\mathop{}\!\mathrm{d}\mu(x)=o(\left\lVert r\right\rVert),$
as $r\to 0$ in $\mathbb{R}^{d}$. For the loss $\ell(t,\vartheta)=2\left\lVert P_{t}-P_{\vartheta}\right\rVert_{\mathsf{TV}}=\left\lVert p_{t}-p_{\vartheta}\right\rVert_{L^{1}}$, we then have $\ell(t,\vartheta)=\left\lVert\langle t-\vartheta,\dot{p}_{\vartheta}\rangle\right\rVert_{L^{1}}+R(t,\vartheta),$ where $R(t,\vartheta)=o(\left\lVert t-\vartheta\right\rVert)$ as $\left\lVert t-\vartheta\right\rVert\to 0$.
Differentiability in quadratic mean, assumed in the BvM theorem (Theorem 1), implies $L^{1}$-differentiability, and moreover, $\dot{p}_{\vartheta}(x)=p_{\vartheta}(x)\ell^{\prime}_{\vartheta}(x)$, where $\ell^{\prime}_{\vartheta}$ is the quadratic mean derivative. Assume that $\vartheta\mapsto\dot{p}_{\vartheta}(x)$ is continuous at $\vartheta_{0}$ for all $x\in\mathcal{X}$, that
$\int_{\mathcal{X}}\sup_{\vartheta\in B_{\delta}(\vartheta_{0})}\left\lVert\dot{p}_{\vartheta}(x)\right\rVert\,\mathop{}\!\mathrm{d}\mu(x)<\infty$
for some $\delta>0$, and that for any compact $K\subset\mathbb{R}^{d}$ and $h\in\mathbb{R}^{d}$,
@@ 5.2.1 Total variation distance
Then condition (iii) in Proposition 3 holds with $\ell_{0}(t,h)=\left\lVert\langle t-h,\dot{p}_{\vartheta_{0}}\rangle\right\rVert_{L^{1}}\text{ and }p=1,$ see Supplement C.5.1. Thus, if the other assumptions of Theorem 2 hold, the Bayes estimator under total variation loss satisfies
$n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})\stackrel{{\scriptstyle D}}{{\longrightarrow}}\arg\min_{t\in\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left\lVert\langle t-h,\dot{p}_{\vartheta_{0}}\rangle\right\rVert_{L^{1}}\,\mathop{}\!\mathrm{d}\mathcal{N}\!\left(Y,I_{\vartheta_{0}}^{-1}\right)(h),$
@@
A direct verification (see Supplement C.5), using $L^{1}$-differentiability and uniform directional differentiability of the Gaussian density, shows that for any compact $K\subset\mathbb{R}^{d}$ and $h\in\mathbb{R}^{d}$,
@@
5.2.2 1-Wasserstein distance
For the particular Stein class $\mathcal{H}=\{f:\mathcal{X}\to\mathbb{R}:\left\lvert f(x)-f(y)\right\rvert\leq\left\lVert x-y\right\rVert\text{ for any }x,y\}$ with compact $\mathcal{X}$, the induced integral probability metric in (7) is the $1$-Wasserstein distance (Villani,, 2003, Theorem 1.14). If $\mathcal{X}$ is one-dimensional (not necessarily compact), the distance admits the distribution function representation (Villani,, 2003, Theorem 2.18): For $P,Q$ with finite first absolute moments and distribution functions $F_{P},F_{Q}$, it holds that $W_{1}(P,Q)=\int_{\mathbb{R}}\left\lvert F_{P}(x)-F_{Q}(x)\right\rvert\,\mathop{}\!\mathrm{d}x=\left\lVert F_{P}-F_{Q}\right\rVert_{L^{1}}.$
@@
If the family is differentiable in quadratic mean, then under mild conditions $F_{\vartheta}$ is $L^{1}$-differentiable with $\dot{F}_{\vartheta}(t)=\int_{-\infty}^{t}\ell^{\prime}_{\vartheta}(x)\,\mathop{}\!\mathrm{d}P_{\vartheta}(x),$ where $\ell^{\prime}_{\vartheta}$ is the quadratic mean derivative (Proposition 6 in Supplement A.2).
@@
Figure 1: Contour plots of maps $(a,b)\mapsto W_{1}(P_{a},P_{b})$ (left) and $(a,b)\mapsto\left\lvert a-b\right\rvert$ (right).
@@ 5.3 Beyond our framework
Here, we cannot employ our framework, since the empirical prior $\Pi_{n}$ is not absolutely continuous with respect to the Lebesgue measure (so that the BvM theorem may fail). Nevertheless, this barycenter can be formulated as the solution of a finite-dimensional linear program, whose standard form (see Supplement C.5.3) is
@@
The barycenter is then expressed as $t=(x_{1},\ldots,x_{d-1},\,1-x_{1}-\cdots-x_{d-1})$. Its limiting distribution (which is non-Gaussian, in general) may be studied via arguments analogous to those in Klatt et al., (2022), who provided limit laws for linear programs with random $b$ in contrast to the random $c$ encountered here. A detailed investigation of this problem is an intriguing direction for future research, but lies beyond the scope of this paper.
@@ 6 Numerical experiments
In this section, we illustrate our theoretical convergence results through numerical simulations. We focus on the three intrinsic losses $\ell_{H},\ell_{\mathrm{KL}},\ell_{W_{2}}$ discussed in Section 5.1.
@@ 6.1 Exponential–Gamma model
$\ell_{H}(t,\vartheta)=1-\frac{2(t\vartheta)^{1/2}}{t+\vartheta},\quad\ell_{\mathrm{KL}}(t,\vartheta)=\log(\vartheta)-\log(t)+\frac{t}{\vartheta}-1,\quad\ell_{W_{2}}(t,\vartheta)=2(t^{-1}-\vartheta^{-1})^{2}.$
In the simulation, we set $a=b=2$ and $\vartheta_{0}=2$. Figure 2 displays QQ-plots of the rescaled and centered estimators $(I_{\vartheta_{0}}n)^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ against the standard normal distribution. As $n$ increases, the sample quantiles approach the theoretical normal quantiles, in agreement with Theorem 3.
Figure 2: $Q$-plots for $(nI_{\vartheta_{0}})^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ for $n\in\{10,50,100,1000\}$ (left to right), based on $M=5000$ Monte Carlo repetitions under the Exponential–Gamma model. Here $\hat{\theta}_{n}$ is the Bayes estimator under the Hellinger loss $\ell_{H}$, the Kullback–Leibler (KL) loss $\ell_{\mathrm{KL}}$ and the 2-Wasserstein loss $\ell_{W_{2}}$ (top to bottom). The reference distribution (shown via theoretical quantiles) is the standard normal distribution.